OpenStudy (anonymous):

1/2-i simplify plz help!

OpenStudy (unklerhaukus):

\[\frac1{2-i}\]?

OpenStudy (unklerhaukus):

\[\frac{1}{2-i}=\frac{1}{2-i}\times1\]\[=\frac{1}{2-i}\times\frac{2+i}{2+i}\]\[=\frac{2+i}{4+i-i-i^2}\]\[=\cdots\]